Question 1 The table below shows Johns Jacks, and its individual demand schedules for apples. Pis the price of apples in U
Answer #1

a Market demand of apples Quantity (Qd) Price (P) 1:00 1:50 2:00 2:50 2:99 (7+7+7 = 21 16+5+6)= 17 115+4+4)= 13 (5+3 + 2) = 1Similarly we can calculate c and d from the calculations of a, b and c